# Approximate Quadratization of High-Order Hamiltonians for Combinatorial Quantum Optimization

### v1

@@ IV Hardware-native circuits for QUBOs
where $E(x)=\langle x|H_{C}|x\rangle$ is the energy of the basis state |$\ket{x}$. The objective value and the energy are related by $2f(x)=|E|-E(x)$. The ground state of $H_{\text{Max-Cut}}$ has $r=1$. On average, uniformly sampled candidate solutions have $r=0.5$. The best known polynomial-time randomized rounding algorithm, the Goemans-Williamson algorithm [62], produces samples with $r\simeq 0.878$, on average. For $s$ shots, an average $r$ is computed over the values $r_{k}$ computed from corresponding bitstrings $x_{k}$.
We generate three random three-regular (RR3) graphs with 40 nodes. Each graph can be SAT mapped following Ref. [53] in $k_{\text{max}}=9$ SWAP layers. Next, we generate an Ansatz circuit with $k\in[0,k_{\text{max}}]$ SWAP layers and depth $p\in\{1,2,3\}$. We train the parameters in each Ansatz using MPS simulators to evaluate the energy with a bond dimension of 20, see App. E.2 for more details. We observe a monotonic increase in the approximation ratio with $k$ evaluated with the MPS, see Fig. 6(a). This shows that the QAOA performance improves as the Ansatz retains more of the problem structure. The circuits are then executed on ibm_fez from which we draw 50 000 noisy samples. The samples at $k=9$ are better than those measured by Sack and Egger on an Eagle device [59], see the black dashed line in Fig. 6(b). This is expected since ibm_fez is a Heron device with lower error rates than Eagle devices. Crucially, this data shows significant gains in solution quality when the truncated Ansatz is applied. In the presence of hardware noise, there is an optimal point $(k,p)$, which represents the maximum depth beyond which increasing either $k$ or $p$ reduces $r$, see Fig. 6(b). On ibm_fez, we find this optimal point at depth-three with two SWAP layers. Once noise begins to reduce the approximation ratio, the higher depth QAOA degrades faster than the lower depth QAOA.
Increasing $k$ and $p$ increases the differences between the ideal and the hardware results, compare Fig. 6(a) and (b). We can recover a noise-free approximation ratio with the ${\rm CVaR}_{\alpha}$ aggregation function [63, 64], which keeps only an $\alpha\in[0,1]$ fraction of the best samples $x$ measured by $f(x)$. Here, we fit $\alpha$ such that the approximation ratio measured with ${\rm CVaR}_{\alpha}$ matches the noiseless one given by the MPS evaluator. The optimization of the alpha parameter is carried out with COBYLA and results in a root mean squared error of $10^{-6}$ between the 90 MPS approximation ratios and the 90 ${\rm CVaR}$ ones (three graphs, three depths and ten values of $k$). As expected, as $k$ and $p$ increase, we discard more shots to recover a noiseless expectation value, see solid lines in Fig. 6(c). This noise strength is then compared to a theoretical amount of noise based on error rates reported by ibm_fez. To this end, we compute $\alpha_{\text{th}}=1/\sqrt{\gamma}$, where $\gamma$ measures the noise strength of the circuit. We make the simple assumption that
@@
Here, $\gamma_{0}$ and $\gamma_{1}$ represent the layer fidelity [65] of the even and odd layer of $CZ$ gates, respectively. The factors $\gamma_{i}^{2p}$ account for the first hardware-native layers of $R_{ZZ}$ gates which transpile to four layers of $CZ$ gates per QAOA layer $p$ on ibm_fez. The terms $\smash{\gamma}_{0}^{3p\lceil\frac{k}{2}\rceil}$ and $\smash{\gamma}_{1}^{3p\lceil\frac{k}{2}\rceil-3p}$ account for the even and odd SWAP layers, respectively, which transpile into three layers of $CZ$ gates each. We approximately express the $\gamma_{i}$’s, where $i\in\{0,1\}$, using the errors $\epsilon_{j,j+1}$ of the native entangling gates between the pairs of qubits $j$ and $j+1$ reported by the backend. For the even and odd layers of two-qubit gates, we have
@@
where the indexing chooses the appropriate two-qubit gate on the line of qubits and we assume $n\geq 3$. This simple $1/\sqrt{\gamma}$ estimation of $\alpha$ already produces excellent qualitative agreement with the fitted $\alpha$, the solid lines in Fig. 6(c) show the same trend as the dashed ones. At a low number of SWAP layers, we overestimate the sampling overhead since the transpiler removes the unnecessary SWAP gates that the SWAP strategy added. When the circuits become deep, e.g. $k=9$ and $p=3$, we tend to underestimate the sampling overhead since we do not account for single-qubit gate noise. In summary, we can (i) trade approximation ratio for a partial implementation of the problem with less noise and (ii) recover a noiseless setting by post-selecting and drawing more samples. This sampling overhead scales exponentially in the circuit depth.
@@ V Discussion and Conclusions
Using noisy quantum hardware to solve non-hardware-native dense optimization problems remains a challenge, which high-order optimization problems exacerbate. For dense four-local Hamiltonians $H_{4}$, the circuit gate count scales as $O(n^{4})$. This scaling limits the size of the LABS problem to about 30 decision variables before the noisy hardware samples are effectively random, for a depth-one QAOA and a depolarizing noise strength of $\lambda=0.001$. This threshold is expected to be even lower for $H_{4}^{\text{full}}$ since it is a denser problem than LABS. We design a scalable Ansatz for high-order optimization problems based on an approximate quadratization of the HUBO using a clique expansion. This construction sacrifices the performance guarantee of QAOA in the $p\to\infty$ limit, but recovers an Ansatz with a depth linear in problem size $n$.
The caveat of our procedure lies in the quadratization method. The clique expansion method, originally designed for mathematical applications [55], does not guarantee that the quadratized Hamiltonian preserves the ground state of the original problem. Indeed, an ideal quatratization would ensure that the ground state of the quadratized Hamiltonian and the quartic one coincide. Our approach is, instead, inherently heuristic, and does not come with a theoretical guarantee.
@@
Our simulations confirm that the standard QAOA Ansatz cannot be used in utility-scale applications which inevitably require error correction. Interestingly, they also show that a quadratized Ansatz yields better samples than the standard QAOA Ansatz once a threshold system size, inversely correlated with the noise strength, is reached. The larger the noise strengths, the faster the quadratized Ansatz starts outperforming the standard one in terms of system size. This makes the quadratized Ansatz particularly appealing for today’s hardware.
@@
Future work may further investigate on which problems this quadratization approach is expected to be maximally efficient. Furthermore, our approach can be extended by developing hybrid methods that mix our approximate quadratization with an exact implementation of a subset of the high-order terms. Given the large number of gates in dense HUBOs, this strategy may require to exactly implement a considerable amount of gates before yielding a relevant improvement over the direct approximate quadratization. This is because the approximate quadratization assigns weights to the resulting graph edges based on the hyperedges it belonged to. If we exactly implement a small number of hyperedges, the resulting graph will consequently only change very little. In addition, one could choose to exactly quadratize the terms in a subset $E^{\prime}$ of the fourth order terms as in [54], which would come at a cost of $2|E^{\prime}|$ qubit overhead.
@@
The second contribution of this paper is designing hardware-friendly Ansatze for QUBOs. We use SWAP Strategies [60, 53, 61, 41] to create a truncated cost Hamiltonian that approximates the target one. We observe that the approximation ratio monotonically increases with the number of SWAP layers $k$ in a noiseless setting. However, this trend changes in the presence of noise, resulting in an optimal pair $(k,p)$ that maximizes the average approximation ratio. In future work, it is worth investigating whether using SWAP Strategies improves the results found in Ref. [41].
@@
Further, optimizing the QAOA variational parameters remains a challenge due to vanishing gradients [66, 8, 67, 68], and an abundance of local minima [69]. Thus, solving dense problems with the typical QAOA Ansatz requires significant quantum and classical resources. Progressively building-up the Ansatz and optimizing its parameters one SWAP layer at a time may mitigate this effect. Future work will investigate the effect of additionally optimizing the weights of the truncated QUBO, similarly to how we quadratized HUBOs.
@@
In summary, our work highlights the challenges in implementing high-order optimization problems on quantum hardware. We show how these challenges can be mitigated by carefully designing the Ansatz circuit from which to sample. Moreover, our work points out additional directions in which possible approximate quadratization may further be explored.
@@ Appendix A Circuit complexity scaling
A.1 Fully-connected QUBO two-qubit gate count
We now give some intuition on the number of entangling gates discussed in Section II and depicted in Fig. 1(a). Assuming all-to-all qubit connectivity, implementing all $n(n-1)/2$ $R_{ZZ}$ terms requires $n(n-1)$ entangling two-qubit gates. Assuming linear qubit connectivity, we use SWAP layers to optimally implement a fully-connected QUBO. A SWAP layer contains $\lfloor n/2\rfloor$ or $\lfloor(n-1)/2\rfloor$ SWAP gates, depending on its parity. This leads to an additional $\Delta N_{\text{line}}=3\times\Big{\lceil}\frac{1}{2}(n-2)\cdot\Big{(}\lfloor n/2\rfloor+\lfloor(n-1)/2\rfloor\Big{)}\Big{\rceil}$ entangling two-qubit gates since a SWAP is equivalent to three CNOT gates. When transpiling the final circuit, the overhead gate count becomes $\Delta N_{\text{line}}/3$ due to the cancellation of neighboring entangling gates, i.e. we have only one entangling gate, instead of three, in the addition to the two corresponding to $R_{ZZ}$ in Fig. 7. The asymptotic factor between circuit gate count on linear and all-to-all connectivity is then
@@ A.1 Fully-connected QUBO two-qubit gate count
where we add the overhead of entangling gates on top of the constant factor $1$, showing that a linear topology requires denser circuits than an all-to-all one. In the $p=1$ case, the fitted data yields $\alpha=1.451$.
Figure 7: Decomposition of an $R_{ZZ}$ gate followed by a SWAP gate into a basis gate set made of single-qubit rotations and a $CX$. Here, two consecutive $CX$ gates cancel. This construction is typical in SWAP networks on linear connectivity, see Fig. 2(d).
A.2 Fully-connected QUBO two-qubit depth
We now rationalize the factor of 1.445 between the circuit depth of a fully-connected QUBO on all-to-all and linear qubit connectivity as seen in Fig. 1(b). A fully-connected QUBO on $n$ variables contains $n(n-1)/2$ terms and is implemented by $n-1$ layers of $R_{ZZ}$ gates with all-to-all connectivity, since each qubit is involved in $n-1$ interactions. Since each $R_{ZZ}$ gate requires two entangling gates, the total two-qubit depth is $2(n-1)$. On a linear qubit connectivity, the same QUBO is optimally implemented using $n-2$ SWAP layers [60]. A SWAP layer has depth three since each SWAP gate is implemented through three entangling gates. In SWAP networks, an $R_{ZZ}$ gate followed by a SWAP gate allows us to simplify to adjacent entangling gates, see Fig. 7. This leads to a total depth of $3(n-2)$ gates. The asymptotic factor between the circuit depth on linear and all-to-all connectivity is then
@@ Appendix B Circuit transpilation
B.1 HUBOs
Here, we describe how we transpile QAOA circuits for the higher-order problems. We rely on the ‘transpile’ function of Qiskit [48], which decomposes physical operators into the basis gate set $\{R_{Z},\sqrt{X},X,CZ\}$. We assume either the default all-to-all connectivity, or a linear one, by providing a coupling map consisting of a list of connected qubit pairs $[i,i+1]$ and $[i+1,i]$, where $i\in[0,n-1]$. We set the optimization level to 3. Qiskit will thus route the QAOA circuit to the hardware topology with the LightSABER algorithm [70] which is an enhancement of the SABER SWAP algorithm [71]. Next, the instructions in the routed circuit will be decomposed to hardware native ones and undergo further optimizations to remove, e.g., canceling gates, and SWAPs as well as diagonal gates that occur before measurement instructions.
B.2 Random three-regular graphs
We now describe the circuit transpilation procedure for the experiments ran on ibm_fez for QAOA for RR3 graphs. First, the qubit connectivity and native gate set are provided by the given backend. We then compile Ansatze circuits with SWAP Strategies as described in the publicly available ‘qopt-best-practices’ repository [72]. For QUBOs that are not fully connected, the circuit depth can often be reduced by appropriately mapping the decision variables onto the physical qubits. We use the approach of Matsuo et al. [53], which finds an initial mapping by solving multiple satisfiability (SAT) problems to minimize the number of SWAP layers. This method can become time-consuming since SAT problems are NP-hard. For our experiments, we set a timeout time of $20$ seconds. After the circuit is built from $R_{ZZ}$ and SWAP gates, we run the Qiskit transpiler on a low optimization level, i.e. 1, to decompose the gates in $CZ$ gates and simplify any redundant gates.
@@ Appendix C Weights of quadratized hypergraphs
In this section, we derive an analytical formula for the weights $w_{2}^{\text{opt}}(j,k)$ of the clique expansion. For now, we consider Hamiltonians which only contain $d$-local terms. Denote by $N_{jk}$ the number of all hyperedges containing edge $(j,k)$. From Eq. (7), we can compactly express $w_{2}^{\text{opt}}(j,k)$. Expanding the binomials and eliminating the constant factors results in
$\arg\min_{w_{2}(j,k)}N_{jk}w_{2}(j,k)^{2}-2w_{2}(j,k)\sum_{e\in\mathcal{E}_{(j,k)}}w_{4}(e),$ (C.1)
@@
$w_{2}^{\text{opt}}(j,k)=\sum_{e\in\mathcal{E}_{(j,k)}}w_{4}(e)\Big{/}N_{jk}.\$ (C.2)
Therefore, the weight of the resulting edges is the mean of the hyperedges it belongs to when the Hamiltonian contains only $d$-local terms. For $H_{4}^{\text{full}}$ and $w_{4}(e)\sim\mathcal{N}(0,1),\forall e\in\mathcal{E}$, we have $w_{2}^{\text{opt}}(j,k)\sim\mathcal{N}(0,N_{jk}^{-1/2}),\forall(j,k)\in E$.
LABS requires a separate analysis since its Hamiltonian also includes two-local terms. We introduce the variable $I_{jk}$ which is $1$ if $Z_{j}Z_{k}$ is in $H_{\text{LABS}}$ and 00 otherwise. We then obtain
$w_{2}^{\text{opt}}(j,k)$ $=\arg\min_{w_{2}(j,k)}$ $\Big{(}I_{jk}(w_{2}(j,k)-1)^{2}$
$+N_{jk}(w_{2}(j,k)-2)^{2}\Big{)}$ (C.3)
$=\arg\min_{w_{2}(j,k)}$ $\Big{(}w_{2}(j,k)^{2}(N_{jk}+I_{jk})$
$-2w_{2}(j,k)(2N_{jk}+I_{jk})\Big{)}$ (C.4)
The step from Eq. (7) to Eq. (C.3) uses the fact that all quartic and quadratic terms have weight 2, and 1, respectively. The edge weights for the quadratized LABS are thus
$w_{2}^{\text{opt}}(j,k)=2-\frac{I_{jk}}{N_{jk}+I_{jk}}.$ (C.5)
@@
The resulting weights $w_{2}^{\text{opt}}(j,k)$ are close to 2 since $I_{jk}\in\{0,1\}$ and $N_{jk}\sim O(n)$. To see why $N_{jk}\sim O(n)$, consider another pair of nodes $(h,l)$ with which the edge $(j,k)$ could form a hyperedge $e=(j,k,h,l)$. The expected value of $N_{jk}$ is the sum of the probabilities that the tuple $(j,k,h,l)$ forms a hyperedge in the original hypergraph $\mathcal{E}$, for all pairs of nodes $h,l\in V$, i.e.
$\mathds{E}(N_{jk})=\sum_{h,l\in V}{\rm Prob}\left[(j,k,h,l)\in\mathcal{E}\right]$ (C.6)

### v2

@@ IV Hardware-native circuits for QUBOs
[…]
We generate three random three-regular (RR3) graphs with 40 nodes. Each graph can be SAT mapped following Ref. [matsuo2023sat] in $k_{\text{max}}=9$ SWAP layers. Next, we generate an Ansatz circuit with $k\in[0,k_{\text{max}}]$ SWAP layers and depth $p\in\{1,2,3\}$. We train the parameters in each Ansatz using an MPS-based simulator to evaluate the energy with a bond dimension of 20, see App. D. We observe a monotonic increase with $k$ of the approximation ratio evaluated with the MPS, see Fig. 6(a). This shows that the QAOA performance improves as the Ansatz retains more of the problem structure. The circuits are then executed on ibm_fez from which we draw 50 000 samples. The samples at $k=9$ are better than those measured by Sack and Egger on an Eagle device [sack2024blackline], see the black dashed line in Fig. 6(b). This is expected since ibm_fez is a Heron device with lower error rates than Eagle devices. Crucially, this data shows significant gains in solution quality when the truncated Ansatz is applied. In the presence of hardware noise, there is an optimal point $(k,p)$, which represents the maximum depth beyond which increasing either $k$ or $p$ reduces $r$, see Fig. 6(b). On ibm_fez, we find this optimal point at depth-three with two SWAP layers. Once noise begins to reduce the approximation ratio, the higher depth QAOA degrades faster than the lower depth QAOA.
Increasing $k$ and $p$ increases the differences between the ideal and the hardware results, compare Fig. 6(a) and (b). We can recover a noise-free approximation ratio with the ${\rm CVaR}_{\alpha}$ aggregation function [barkoutsos2020cvar, barron2024cvar], which keeps only an $\alpha\in[0,1]$ fraction of the best samples $x$ measured by $f(x)$. Here, we fit $\alpha$ such that the approximation ratio measured with ${\rm CVaR}_{\alpha}$ matches the noiseless one obtained with the MPS simulator. The optimization of the $\alpha$ parameter is carried out with COBYLA and results in a root mean squared error of $10^{-6}$ between the 90 MPS approximation ratios and the 90 $\rm CVaR$} ones (three graphs, three depths and ten values of $k$). As expected, as $k$ and $p$ increase, we discard more shots to recover a noiseless expectation value, see solid lines in Fig. 6(c). This noise strength is then compared to a theoretical amount of noise based on error rates reported by ibm_fez. To this end, we compute $\alpha_{\text{th}}=1/\sqrt{\gamma}$, where $\gamma$ measures the noise strength of the circuit. We make the simple assumption that
@@
Here, $\gamma_{0}$ and $\gamma_{1}$ represent the layer fidelity [mckay2023benchmarking] of the even and odd layer of $CZ$ gates, respectively. The factors $\gamma_{i}^{2p}$ account for the first hardware-native layers of $R_{ZZ}$ gates which transpile to four layers of $CZ$ gates per QAOA layer $p$ on ibm_fez. The terms $\smash{\gamma}_{0}^{3p\lceil\frac{k}{2}\rceil}$ and $\smash{\gamma}_{1}^{3p\lceil\frac{k}{2}\rceil-3p}$ account for the even and odd SWAP layers, respectively, which transpile into three layers of $CZ$ gates each. We approximately express $\gamma_{i}$, where $i\in\{0,1\}$, using the errors $\epsilon_{j,j+1}$ of the native entangling gates between the pairs of qubits $j$ and $j+1$ reported for the used quantum processor. For the even and odd layers of two-qubit gates, we have
@@
where the indexing is based on an appropriate indexing of the two qubit gates and we assume $n\geq 3$. This simple $1/\sqrt{\gamma}$ estimation of $\alpha$ already produces excellent qualitative agreement with the fitted $\alpha$, the solid lines in Fig. 6(c) show the same trend as the dashed ones. At a low number of SWAP layers, we overestimate the sampling overhead since the transpiler removes the unnecessary SWAP gates that the SWAP strategy added. When the circuits become deep, e.g. $k=9$ and $p=3$, we tend to underestimate the sampling overhead since we do not account for single-qubit gate noise. In summary, we can (i) trade approximation ratio for a partial implementation of the problem with less noise and (ii) recover a noiseless setting by post-selecting and drawing more samples. This sampling overhead scales exponentially in the circuit depth.
@@ V Discussion and Conclusions
Using noisy quantum hardware to solve non-hardware-native dense optimization problems remains a challenge, which high-order optimization problems exacerbate. For dense four-local Hamiltonians, the circuit gate count scales as $O(n^{4})$. The circuit depth limits the size of the LABS problem to about 30 decision variables before the noisy hardware samples are effectively random, for a depth-one QAOA and a depolarizing noise strength of $\lambda=0.001$. This threshold is expected to be even lower for denser problems than LABS.
To reduce the circuit depth without increasing its width we explore approximate quadratizations of the HUBO problem. These constructions sacrifice the performance guarantee of QAOA in the $p\to\infty$ limit, but recover an Ansatz with a depth linear in problem size $n$. The clique expansion method, originally designed for mathematical applications [suppakitpaisarn2024utilizing], does not guarantee that the quadratized Hamiltonian preserves the ground state of the original problem. Indeed, an ideal quatratization would ensure that the ground state of the quadratized Hamiltonian and the quartic one coincide. By contrast, optimizing the parameters of a quadratized Hamiltonian Ansatz partially overcomes this issue. The resulting samples have a high probability to have an approximation ratio close to one while being sub-optimal.
@@
Our simulations confirm that the standard QAOA Ansatz cannot be used in utility-scale applications because it inevitably requires error correction. Interestingly, they also show that a quadratized Ansatz yields better samples than the standard QAOA Ansatz once a threshold noise strength is reached. This makes the quadratized Ansatz particularly appealing for current quantum hardware.
@@
Future work may further investigate on which problems this quadratization approach may be maximally efficient. Furthermore, our approach can be extended by developing hybrid methods that mix our approximate quadratization with an exact implementation of a subset of the high-order terms. In addition, one could choose to exactly quadratize the terms in a subset $E^{\prime}$ of the fourth order terms as in [mandal2020compressed], which would come at a cost of $2|E^{\prime}|$ qubit overhead.
@@
The second contribution of this paper is designing hardware-friendly Ansatze for QUBOs. We use SWAP Strategies [weidenfeller2202scaling, matsuo2023sat, Harrigan_2021, maciejewski2024design] to create a truncated cost Hamiltonian that approximates the target one. The resulting approximation ratio monotonically increases with the number of SWAP layers $k$ in a noiseless setting. However, in the presence of noise, there is an optimal pair $(k,p)$ that maximizes the average approximation ratio. In future work, it is worth investigating whether using SWAP Strategies improves the results found in Ref. [maciejewski2024design].
@@
Further, optimizing the QAOA variational parameters remains a challenge due to vanishing gradients [larocca2024review, holmes2022connecting, arrasmith2022equivalence, mcclean2018barren], and an abundance of local minima [bittel2021training]. Thus, solving dense problems with the typical QAOA Ansatz requires significant quantum and classical resources. Progressively building-up the Ansatz and optimizing its parameters one SWAP layer at a time may mitigate this effect. Future work will investigate the effect of additionally optimizing the weights of the truncated QUBO, similarly to our HUBOs quadratization scheme.
@@
In summary, our work highlights the challenges in implementing high-order optimization problems on quantum hardware. We show how these challenges can be mitigated by carefully designing the Ansatz circuit from which to sample.
@@ Appendix A Circuit complexity scaling
We now give some intuition on the number of entangling gates discussed in Section II and depicted in Fig. 1(a). Assuming all-to-all qubit connectivity, implementing all n(n1)/2n(n-1)/2 $R_{ZZ}$ terms requires n(n1)n(n-1) entangling two-qubit gates with a depth of 2(n1)2(n-1). Assuming linear qubit connectivity, we use n2n-2 SWAP layers to optimally implement a fully-connected QUBO. A SWAP layer contains $\lfloor n/2\rfloor$ or $\lfloor(n-1)/2\rfloor$ SWAP gates, depending on its parity. This leads to an additional $\Delta N_{\text{line}}=3\times\Big\lceil\frac{1}{2}(n-2)\cdot\Big(\lfloor n/2\rfloor+\lfloor(n-1)/2\rfloor\Big)\Big\rceil$ entangling two-qubit gates since a SWAP is equivalent to three CNOT gates. When transpiling the final circuit, the overhead gate count becomes $\Delta N_{\text{line}}/3$ due to the cancellation of neighboring entangling gates, i.e. we have only one entangling gate, instead of three, in the addition to the two corresponding to $R_{ZZ}$ in Fig. 7. The asymptotic factor between circuit gate count on linear and all-to-all connectivity is then
@@
where we add the overhead of entangling gates on top of the constant factor $1$. The asymptotic factor between the circuit depth on linear and all-to-all connectivity is then
@@
Figure 7: Decomposition of an $R_{ZZ}$ gate followed by a SWAP gate into a basis gate set made of single-qubit rotations and a CXCX. Here, two consecutive CXCX gates cancel. This construction is typical in SWAP networks on linear connectivity, see Fig. 2(d).
@@ Appendix B Circuit transpilation
Here, we describe how we transpile QAOA circuits for the higher-order problems. We rely on the ‘transpile’ function of Qiskit [javadi2024quantum], which decomposes physical operators into the basis gate set $\{R_{Z},\sqrt{X},X,CZ\}$. We assume either the default all-to-all connectivity, or a linear one, by providing a coupling map consisting of a list of connected qubit pairs $[i,i+1]$ and $[i+1,i]$, where $i\in[0,n-1]$. We set the optimization level to 3. Qiskit will thus route the QAOA circuit to the hardware topology with the LightSABER algorithm [zou2024lightsabre] which is an enhancement of the SABER SWAP algorithm [li2019mapping]. Next, the instructions in the routed circuit will be decomposed to hardware native ones and undergo further optimizations to remove, e.g., canceling gates, and SWAPs as well as diagonal gates that occur before measurement instructions.
We now describe the circuit transpilation for the experiments ran on ibm_fez for QAOA for RR3 graphs. First, the qubit connectivity and native gate set are provided by the given backend. We then compile Ansatze circuits with SWAP Strategies as described in the publicly available ‘qopt-best-practices’ repository [SWAPStrategies_code]. For QUBOs that are not fully connected, the circuit depth can often be reduced by appropriately mapping the decision variables onto the physical qubits. We use the approach of Matsuo et al. [matsuo2023sat], which finds an initial mapping by solving multiple satisfiability (SAT) problems to minimize the number of SWAP layers. This method can become time-consuming since SAT problems are NP-hard. For our experiments, we set a timeout time of $20$ seconds. After the circuit is built from $R_{ZZ}$ and SWAP gates, we run the Qiskit transpiler on a low optimization level, i.e. 1, to decompose the gates in $CZ$ gates and simplify any redundant gates.
@@ Appendix C Weights of quadratized hypergraphs
In this section, we derive an analytical formula for the weights $w_{ij}$ of the clique expansion. From Eq. (7), we can compactly express $w_{ij}$. Expanding the binomials and eliminating the constant factors results in
$\arg\min_{w_{ij}}\sum_{e\in\mathcal{E}:(i,j)\in E(e)}\alpha_{e}w_{ij}^{2}-2w_{ij}\alpha_{e}w_{e},$ (C.1)
@@
$w_{ij}=\sum_{e\in\mathcal{E}:(i,j)\in E(e)}w^{2}_{e}\Big/\sum_{e\in\mathcal{E}:(i,j)\in E(e)}w_{e}.\\$ (C.2)
Therefore, the weight of the resulting edges is a weighted mean of the hyperedges it belongs to.
The weights of the quadratization resulting from LABS can be more precisely calculated using the fact that all quartic and quadratic terms have weight 2, and 1, respectively. We introduce the variable $I_{ij}$ which is $1$ if $Z_{i}Z_{j}$ is in $H_{\text{LABS}}$ and 0 otherwise. We also denote by $N_{ij}$ the number of all hyperedges containing edge $(i,j)$. We then obtain from Eq. (7)
$w_{ij}$ $=\arg\min_{w_{ij}}$ $\Big(I_{ij}(w_{ij}-1)^{2}$
$+\sum_{e\in\mathcal{E}:(i,j)\in E(e)}2(w_{ij}-2)^{2}\Big)$ (C.3)
$=\arg\min_{w_{ij}}$ $\Big(w_{ij}^{2}(2N_{ij}+I_{ij})$
$-2w_{ij}(4N_{ij}+I_{ij})\Big)$ (C.4)
The edge weights for the quadratized LABS are thus
$w_{ij}=2-\frac{I_{ij}}{2N_{ij}+I_{ij}}.$ (C.5)
@@
The resulting weights $w_{ij}$ are close to 2 since $I_{ij}\in\{0,1\}$ and $N_{ij}\sim O(n)$. To see why $N_{ij}\sim O(n)$, consider another pair of nodes $(h,l)$ with which the edge $(i,j)$ could form a hyperedge $e=(i,j,h,l)$. The expected value of $N_{ij}$ is the sum of the probabilities that the tuple $(i,j,h,l)$ forms a hyperedge in the original hypergraph $\mathcal{E}$, for all pairs of nodes $h,l\in V$, i.e.
$\mathds{E}(N_{ij})=\sum_{h,l\in V}{\rm Prob}\left[(i,j,h,l)\in\mathcal{E}\right]$ (C.6)